OpenStudy (anonymous):

In parallelogram ABCD, mB = 8x - 25 and mC = 7x + 55. mB = ?

OpenStudy (anonymous):

use <B + <C = 180

OpenStudy (anonymous):

8x - 25 + 7x + 55 = 180

OpenStudy (anonymous):

you add them together and set it equal to 180

OpenStudy (anonymous):

15x = 150 so x= 10

OpenStudy (anonymous):

mB = 8x - 25=80-25 =55